# Topological piezomagnetic effect in two-dimensional Dirac quadrupole altermagnets

### v1

@@ Appendix D Topological transition in the Lieb lattice altermagnet
[…]
The expansion of the Lieb lattice Hamiltonian around the $X$ and $Y$ points follows the same recipe as the expansion described in Appendix B. Expanding the function $t_{x,{\bf k}}$ yields
@@
To proceed, we write $N_{z}$ as $N_{z}=N_{z,c}+m$, such that $\delta$ can be viewed as a band inversion parameter; $m<0$ describes the topological (mirror Chern insulator) phase, whereas $m>0$ describes the trivial phase. With $N_{z}$ written in this way, the linearized continuum model Hamiltonian at $X$ and $Y$ takes the form
@@
From this result it follows that at $m=0$ the polarizability jumps by an amount $4t_{0}/\pi t_{1}$, in units of $\mu_{B}/a^{2}$. This is precisely what we find in Fig. 4: the discontinuity at $m=0$ found in the full lattice calculation is in perfect agreement with the result from the Dirac model calculationm. Furthermore, as mentioned, in the continuum model calculation only the geometric term contributes to the polarizability, which is confirmed by the numerical lattice model calculation shown in Fig. 4(b). In Fig. 4(b) we separately show the contributions from the geometric and the interband terms, revealing that only the geometric term exhibits a discontinuity at the topological transition.

### v2

@@ Appendix F Topological transition in the Lieb lattice altermagnet
In Fig. 4(a) we show the polarizability $\lambda$ as a function of the magnetic order parameter $N_{z}$, in units of $N_{z,c}$. A key feature is the discontinuity of the polarizability right at the topological transition, when $N_{z}=N_{z,c}$ (here we choose $N_{z}>0$ for concreteness). At this topological transition the gap closes and the system becomes a trivial insulator with vanishing mirror Chern number. The gap closing transition is described by a Dirac fermion mass inversion at both $X=(\pi,0)$ and $Y=(0,\pi)$. The discontinuity of $\Lambda$ can be captured by a linearized continuum model obtained from expanding the lattice Hamiltonian in small momentum $\bf q$} around the $X$ and $Y$ points. Since here we focus on $N_{z}>0$ for concreteness, it is straightforward to determine that the gap closing transition at $X$ ($Y$) is in the $\sigma=\downarrow$ ($\sigma=\uparrow$) sector.
The expansion of the Lieb lattice Hamiltonian around the $X$ and $Y$ points follows the same recipe as the expansion described in Appendix D. Expanding the function $t_{x,{\bf k}}$ yields
@@
To proceed, we write $N_{z}$ as $N_{z}=N_{z,c}+m$, such that $m$ can be viewed as a band inversion parameter; $m<0$ describes the topological (mirror Chern insulator) phase, whereas $m>0$ describes the trivial phase. With $N_{z}$ written in this way, the linearized continuum model Hamiltonian at $X$ and $Y$ takes the form
@@
From this result it follows that at $m=0$ the polarizability jumps by an amount $4t_{0}/\pi t_{1}$, in units of $\mu_{B}/a^{2}$. This is precisely what we find in Fig. 4: the discontinuity at $m=0$ found in the full lattice calculation is in perfect agreement with the result from the Dirac model calculation. Furthermore, as mentioned, in the continuum model calculation only the geometric term contributes to the polarizability, which is confirmed by the numerical lattice model calculation shown in Fig. 4(b). In Fig. 4(b) we separately show the contributions from the geometric and the interband terms, revealing that only the geometric term exhibits a discontinuity at the topological transition.